OpenStudy (anonymous):

how do you divide radicals with integers? like 4√2 divided by √3. does it just equal 4√2 over 3 in fraction form to signify division? or does it have a real answer?

OpenStudy (anonymous):

oops, i meant 4√2 over √3, not over 3

OpenStudy (saifoo.khan):

\[\frac{4\sqrt2}{\sqrt3}\] \[\frac{4\sqrt2}{\sqrt3} \times \frac{\sqrt3}{\sqrt3}\] \[\frac{4\sqrt6}{3}\]

OpenStudy (anonymous):

alright thanks i see. just curious, can you simplify further with the √6? like can you get √6 / 3 and get √2?

OpenStudy (saifoo.khan):

o_O please rewrite your statement..

OpenStudy (anonymous):

lol sorry. like can you divide √6 by 3 and get √2? is it possible to divide radicals with regular integers?

OpenStudy (saifoo.khan):

Nope, that's the final answer.

OpenStudy (saifoo.khan):

If u had, √6 by √3 then it will be √2

OpenStudy (anonymous):

oh alright thanks. that answers my other question i just posted too lol, thank you

OpenStudy (saifoo.khan):

lol, Welcome.